OpenStudy (anonymous):

The probability that james will win two races in the next two days is 1/8 for tomorrow and 1/10 for the day after tomorrow. What is P(win tomorrow, then win the after tomorrow)?

OpenStudy (anonymous):

P(Two wins today and tomorrow) = P(win today)*P(win tomorrow) = (1/8)(1/10) = (1/80)

OpenStudy (anonymous):

thankyou